OpenStudy (anonymous):

help me please!! write the equation of each parabola in vertex form. the two points i am given is (0,9) & (2,1)

OpenStudy (anonymous):

here is the graph

OpenStudy (mathmale):

The very first thing I notice in your graph is the the point (2,1) represents the vertex of the parabola, and the other point, (0,9), the vertical intercept. The simplest form of the equation of a vertical parabola is y=x^2. This parabola has its vertex at (0,0). If we move the whole graph 2 units to the right, the vertex is now at (2,0), but the shape is exactly the same as before. The equation is y=a(x-2)^2. If we move the whole graph both 2 units to the right and 1 unit up, the vertex is now at (2,1), but the shape is exactly the same as before. The equation is y-1=a(x-2)^2. This is our immediate situation. We know that the graph goes through (0,9). Substitute x=0 and y=9 into the most recent equation, above, determine the value of a, and then write the equation again, including that "a" value.

OpenStudy (anonymous):

ok so would the equation be 9-1=a(0-2)^2

OpenStudy (anonymous):

and would it be a=2

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Saige: 9-1=a(0-2)^2 simplifies to 8=a(-2)^2. Mind solving that for a?

OpenStudy (anonymous):

well it would be a4 so wouldn't you divide both sides by 4 to get a=2

OpenStudy (mathmale):

What I'd like to do here is to ask you to check whether this a=2 is correct or not. First, write out the solution you've already proposed, with a = 2: y=2(x-2)^2+1. Now substitute x=0. You should get 9. Do you? Now substitute x = 2. You should get 1. Do you? You should know by then whether or not your a =2 is correct. Good going.

OpenStudy (anonymous):

yes you get 9 ans 1. so a=2 is correct

OpenStudy (mathmale):

Perfect. Congrats. Any questions about this process?

OpenStudy (anonymous):

no i get that part know i just dont get how to put it all i vertex form

OpenStudy (mathmale):

In general, the equation of a parabola with vertex at (a,b) is y=a(x-1) + b. Some people might write it as y-b=a(x-a)^2. I thought it'd be a bit more helpful to show you how moving the vertex affects the equation y=ax^2.

OpenStudy (mathmale):

Please explain in a bit more detail what you mean by "put it all in vertex form."

OpenStudy (anonymous):

i need to put the graph that i attached in vertex form. i do not know how

OpenStudy (mathmale):

Oh, I see. Actually, we've already done that. What is the vertex as shown in the graph? What is the one other point on that graph?

OpenStudy (anonymous):

the vertex was (2,1) and the other point was (0,9)

OpenStudy (anonymous):

so is the vertex form the equation y=2(x-2)^2+1.

OpenStudy (mathmale):

That's right. As I mentioned a moment back, the general form for the equation of a parabola with vertex at (a,b) is y=a(x-a)^2 + b. You correctly identified the v ertex from the graph, so you know that a=2 and b=1. Substitute those two values into the general form. What do you get?

OpenStudy (mathmale):

Excuse me, you've already done everything you needed to do. How satisfied are you that you understand what to do next time?

OpenStudy (anonymous):

i understand how to do it all now. thank you so much. thanks for your time.

OpenStudy (mathmale):

My great pleasure, Saige! See you.